# Compositional simulation-based inference for time series

### v1

@@ A Mixture of Gaussian Random Walk of form
The periodic SDE is a linear SDE parameterized as follows
@@
which gives us a stochastic oscillator. Each transition corresponds to simulating this SDE for $0.1$ ms, using 20 Euler Maruyama steps. As proposal, we choose $\tilde{p}({\bm{x}}^{t})=\mathcal{N}({\bm{x}}^{t};\bm{0},{\bm{I}})$. For evaluation, we fixed the initial distribution to $p({\bm{x}}^{0})=(-0.5,0.5)^{T}$, to quickly start with oscillations.
@@
where every entry of ${\bm{A}}$ and ${\bm{B}}$ is directly parameterized by ${\bm{\theta}}\in\mathbb{R}^{18}$ and ${\bm{x}}^{t}\in\mathbb{R}^{3}$. Reference posterior samples are obtained in the same way as for the previous task.
@@
which samples from a double-well potential with modes position depending on $\theta_{1},\theta_{2}$ (Singer, 2002). We use a combination of particle filter (Doucet et al., 2009) and pseudo-marginal-like Metropolis-Hastings MCMC (Andrieu & Roberts, 2009) sampler to generate reference posterior samples.
@@ B.2 Training and evaluation
For FNLE and FNRE we use adapted reference implementation as implemented in the sbi package (Tejero-Cantero et al., 2020b). In FNLE, we use a 5-layer Masked Autoregressive Flow (Papamakarios et al., 2017), each parameterized by a 2-layer MLP with a hidden dimension of 50. In FNRE we use resnet classifier with two blocks each considering 2 layers with hidden dimension of 50. As MCMC sampling algorithm we use a per-axis slice sampling algorithm. To avoid mode-collapse, we run 100 parallel chains.
@@
[…]
@@
We use an AdamW optimizer with a learning rate of $45e-$ 4 with a cosine schedule and a training batch size of $1000$. Similar to the SBI routine, we use early stopping, but with a maximum number of $5000$ epochs.
Appendix C Additional Experiments
C.1 Time-Variant Gaussian Random Work
To evaluate the performance of the time-variant extension described in Appendix A.2.1, we consider the Time-Variant Gaussian Random Walk (TV Gaussian RW)
with ${\bm{x}}_{t}$ multiplied by the time-variant coefficient $\left(\alpha+\frac{1}{t+1}\right)$. We demonstrate the scalability of our methods with dimension $d$ ranging from $1,2$, to $10$, and the number of observation transitions ranging from $1$, $10$, to $100$. As a comparison metric, we use C2ST. The simulation budget is fixed at $10k$ and $100k$, consistent with the problem settings in the main section. Note that the training data include the time variable $t$, and therefore, the training data for estimating the transition distribution $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$ are reduced to $\frac{1}{100}$ compared to the time-invariant setting in the main section. The maximum time length $T_{\text{max}}$ in training is set to $100$.
C.2 Score composition rules
We utilize the GAUSS composition rule as our default method:
$\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a}|{\bm{x}}^{0:T})\approx\mathbf{\Lambda}({\bm{\theta}}_{a})^{-1}\left(\sum_{t=0}^{T-1}\mathbf{\Sigma}_{a,t,t+1}^{-1}s_{\phi}({\bm{\theta}}_{a}|{\bm{x}}^{t},{\bm{x}}^{t+1})+(1-T)\mathbf{\Sigma}_{a}^{-1}\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a})\right)$
with
$\mathbf{\Lambda}({\bm{\theta}}_{a})=\sum_{t=0}^{T-1}\mathbf{\Sigma}_{a,t,t+1}^{-1}+(1-T)\mathbf{\Sigma}_{a}^{-1}.$
We estimate the posterior precision matrices $\mathbf{\Sigma}_{a,t,t+1}^{-1}$ as recommended by Linhart et al. (2024). For each posterior $p({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$, we produce $500\cdot\dim_{\bm{\theta}}$ samples in parallel from the diffusion model. These samples are then used to estimate the associated covariance matrix. Although this approach introduces a slight constant overhead prior to each sampling run, it does not require the iterative estimation of Jacobians, as is necessary with the JAC method.
Importantly, since the GAUSS method is derived under Gaussian assumptions, it presupposes that the posterior covariance matrices (and their corresponding scores) are smaller than those of the prior due to observational constraints. This assumption may not hold true, particularly in multi-modal tasks. For instance, in the Periodic SDE task, four modes are symmetrically positioned around the origin. While the variances of each mode may decrease, the total variance often still increases, especially when the modes are distant from the origin. This situation can render $\Lambda({\bm{\theta}}_{a})$ non-positive definite, thus invalidating the approximation.
Overall, this strategy significantly enhances the performance of both GAUSS and JAC, even when their underlying assumptions are not fully validated (see Fig. 7 for benchmarks). This improvement is particularly critical for scaling to long time series, especially when the number of function evaluations is reduced (see Fig. 7b).
@@ C.2 Score composition rules
It is noteworthy that the GAUSS and JAC methods yield theoretically equivalent results in the Gaussian RW task, with no violations of assumptions (the only difference lies in their approach to estimating posterior precision matrices). As such, the performance of both methods is quite similar (Fig. 7, first column). However, when assumptions are violated, the differences between the two approaches become apparent. The GAUSS method presumes a Gaussian posterior and analytically computes the associated denoising posteriors. In contrast, the JAC method iteratively derives the best Gaussian approximation for the denoising posterior, as discussed by Boys et al. (2024). Notably, for samplers with fewer steps on non-Gaussian tasks, JAC tends to perform slightly better than GAUSS on long time series (Fig. 7b, columns two onwards). The imposed regularity conditions appear to resolve the numerical challenges associated with this approach, as observed by Linhart et al. (2024).
C.3 Proposal distribution
@@
The choice of proposal distribution can either be an advantage or a challenge. In our benchmark tasks, we primarily selected simple yet reasonable distributions based on our prior knowledge of the dynamics involved. However, we examine the impact of this choice more closely in the Gaussian RW and Mixture RW tasks (see Fig. 8).
@@
We set the proposal distribution to a Gaussian with a zero mean and variable variance, ranging from very narrow to quite wide. Importantly, at evaluation, we initialize the distribution as a point mass at zero. For a very narrow proposal, we anticipate strong performance at $T=1$. However, as $T$ increases, performance may decline because the time series dynamics could extend beyond the high-probability support of the proposal. Conversely, excessively wide proposals may also hinder performance if the values at the observation time fall outside the training set. It is crucial to note that this performance evaluation is relative to our specific evaluation pipeline. In both tasks, evaluations are conducted on random walks that consistently start from zero, thus constraining the range of attainable values given a maximum number of steps.
@@
It is important to emphasize that the proposals we utilized are generally not optimal. Other tasks, such as the Periodic/Linear SDE, Double Well, SIR, and Lotka-Volterra models, exhibit a high correlation among different variables, which the proposal we chose does not capture. We believe that identifying a more suitable proposal could greatly enhance the efficiency of our methods, although doing so may prove nontrivial (and might, itself, require many simulation calls). Nevertheless, to ensure a fair comparison with the NPE baseline, we opted for relatively simple proposals.
C.4 Kolmogorov flow
@@
It is important to note that in the NPE approach, we would need to construct an embedding network not only for individual ”images” but also for a complete video of arbitrary length, which poses significant challenges.
As an initial distribution, we utilize a filtered velocity field provided by jax_cfg, with a maximum velocity and peak wave number of three. Our approach closely follows the methodology outlined by Rozet & Louppe (2023). The simulator performs 100 solve steps per transition, with a step size of $10^{-3}$ seconds; each transition thus emulates 10 milliseconds. To introduce stochasticity into the PDE, we add small amounts of Gaussian noise with a standard deviation of $5\cdot 10^{-3}$ after each transition. The simulations are conducted on a regular $64\times 64$ grid.
@@
Notably, this procedure does not violate the requirement that $\tilde{p}({\bm{x}}^{T})$ is independent of ${\bm{\theta}}$, as each transition is performed with a different, independent parameter.
@@
Fig 10: Time Variant Gaussian random walk: We train all methods on 10k (top row) and 100k (bottom row) simulation on the time-variant Gaussian RW talk. All methods have also been adapted to input the current time (as given by integers ranging from zero to 100).

### v2

@@ A Mixture of Gaussian Random Walk of form
The periodic SDE is a linear SDE parameterized as follows:
@@
which gives us a stochastic oscillator. Each transition corresponds to simulating this SDE for $0.1$ ms, using 20 Euler Maruyama steps. As proposal, we choose $\tilde{p}({\bm{x}}^{t})=\mathcal{N}({\bm{x}}^{t};\bm{0},{\bm{I}})$. For evaluation, we fixed the initial distribution to $p({\bm{x}}^{0})=\delta((-0.5,0.5)^{T})$, to quickly start with oscillations.
@@
where every entry of ${\bm{A}}$ and ${\bm{B}}$ is directly parameterized by ${\bm{\theta}}\in\mathbb{R}^{18}$ and ${\bm{x}}^{t}\in\mathbb{R}^{3}$. Reference posterior samples are obtained in the same way as for the previous task. As proposal we choose $\tilde{p}({\bm{x}}^{t})=\mathcal{N}({\bm{x}}^{t},\mathbf{0},\mathbf{I})$. For evaluation, we fixed the initial distribution to $p({\bm{x}}^{0})=\delta(\mathbf{0})$.
@@
which samples from a double-well potential with modes position depending on $\theta_{1},\theta_{2}$ (Singer, 2002). As proposal we use $p({\bm{x}}^{t})=\text{Unif}(-2.5,2.5)$, as intial distirbution we at evaluation we use $p(x_{0})=\delta((1,-1,1,-1)^{T})$ We use a combination of particle filter (Doucet et al., 2009) and pseudo-marginal-like Metropolis-Hastings MCMC (Andrieu & Roberts, 2009) sampler to generate reference posterior samples.
@@ D.2 Training and evaluation
For FNLE and FNRE, we use adapted reference implementation as implemented in the sbi package (Tejero-Cantero et al., 2020b). In FNLE, we use a 5-layer Masked Autoregressive Flow (Papamakarios et al., 2017), each parameterized by a 2-layer MLP with a hidden dimension of 50. In FNRE, we use resnet classifier with two blocks each considering 2 layers with hidden dimension of 50. As MCMC sampling algorithm, we use a per-axis slice sampling algorithm. To avoid mode-collapse, we run 100 parallel chains.
@@
We set $\beta_{\text{min}}=0.1$, and $\beta_{\text{max}}=10$ for all experiments. Both for the time interval $[10^{-2},1.]$. The associated conditional means and variances can be derived from this SDE (Song et al., 2021).
@@
We use an AdamW optimizer with a learning rate of $5\cdot 10^{-4}$ with a cosine schedule and a training batch size of $1000$. Similar to the SBI routine, we use early stopping, but with a maximum number of $5000$ epochs.
Appendix E Additional Experiments
E.1 Extended experimental comparison
In this section, we extended the empirical evaluation and collected the results in Table 3 and 4. We additionally added the following baseline comparisons:
  • NPE’ has an alternative embedding network (no RNN). Inspired by the fact that our factorized methods work on all windows (${\bm{x}}^{t}$, ${\bm{x}}^{t+1}$), we embed all such windows into 200 dim feature vectors (using an MLP). All feature vectors are accumulated (using a summation) and mapped into a final 50-dim summary statistics (using another MLP). This is also similar to the embedding networks used in the i.i.d. case (Zaheer et al., 2018).
  • NPE50 corresponds to our standard NPE with RNN embedding net. By default, we do train NPE on sequences of length ten; here, we extended this to $50$ (although this means that the number of joint samples decreases accordingly, i.e., for 10k, only 200 for 100k 2000 to ensure an overall equivalent training budget).
  • NPE+ with pre-trained neural sufficient summary statistic. We follow (Chen et al., 2021) infomax learning approach using the distance correlation objective to pre-train the embedding net.
  • NPE++ with pre-trained sliced neural sufficient summary statistic (Chen et al., 2023).
  • NLE+: NLE with neural sufficient summary statistic (Chen et al., 2021).
  • NLE++: NLE with neural sliced sufficient summary statistic (Chen et al., 2023).
  • NSE Neural posterior score estimation (Sharrock et al., 2024; Geffner et al., 2023)(analogous to our standard NPE baseline, but with score estimation).
Task Periodic SDE Linear SDE Mixture RW Double Well
T=1 T=10 T=100 T=1 T=10 T=100 T=1 T=10 T=100 T=1 T=10 T=100
NPE 0.82 0.84 0.96 0.72 0.93 0.97 0.92 0.99 1.00 0.64 0.67 0.81
NPE’ 0.78 0.91 0.99 0.70 0.93 0.99 0.91 0.99 1.00 0.55 0.72 0.99
NPE50 0.80 0.93 0.98 0.75 0.93 0.97 0.92 0.99 1.00 0.58 0.68 0.85
NPE+ 0.80 0.89 0.99 0.75 0.92 0.97 0.91 0.99 1.00 0.67 0.65 0.91
NPE++ 0.79 0.90 0.97 0.74 0.92 0.97 0.91 0.99 1.00 0.64 0.64 0.83
NLE+ 0.80 0.88 0.97 0.76 0.92 0.97 0.96 0.99 1.00 0.69 0.64 0.78
NLE++ 0.75 0.93 0.98 0.76 0.93 0.97 0.95 0.99 1.00 0.55 0.60 0.81
NSE 0.77 0.92 0.97 0.68 0.93 0.97 0.91 0.99 1.00 0.60 0.59 0.79
FNLE 0.57 0.66 0.83 0.57 0.78 0.98 0.71 0.89 0.98 0.52 0.64 0.87
FNRE 0.63 0.89 0.98 0.69 0.92 0.99 0.91 0.99 1.00 0.64 0.89 0.97
FNSE 0.56 0.65 0.85 0.56 0.82 0.97 0.57 0.78 0.92 0.52 0.68 0.89
NPE 0.68 0.63 0.95 0.70 0.84 0.96 0.91 0.99 1.00 0.62 0.64 0.76
NPE’ 0.62 0.62 0.99 0.67 0.74 1.00 0.91 0.99 1.00 0.62 0.62 0.99
NPE50 0.81 0.83 0.87 0.71 0.92 0.97 0.91 0.99 1.00 0.63 0.69 0.76
NPE+ 0.72 0.69 0.99 0.66 0.89 0.96 0.91 0.99 1.00 0.60 0.60 0.80
NPE++ 0.67 0.65 0.95 0.71 0.91 0.97 0.91 0.99 1.00 0.65 0.61 0.78
NLE+ 0.67 0.82 0.97 0.71 0.91 0.97 0.91 0.99 1.00 0.55 0.58 0.81
NLE++ 0.73 0.8 0.96 0.70 0.92 0.97 0.92 0.99 1.00 0.55 0.58 0.79
NSE 0.71 0.65 0.96 0.66 0.77 0.93 0.91 0.99 1.00 0.54 0.54 0.77
FNLE 0.54 0.60 0.72 0.53 0.69 0.89 0.61 0.76 0.89 0.50 0.54 0.75
FNRE 0.55 0.75 0.92 0.54 0.81 0.99 0.70 0.86 0.93 0.50 0.74 0.90
FNSE 0.53 0.60 0.76 0.53 0.70 0.90 0.52 0.69 0.69 0.50 0.59 0.81
Table 3: Extended baseline comparison: Performance per method given as C2ST metric for each benchmark task. The top half of the table reports results trained on 10k step simulations, and the bottom half on 100k. Best performing methods are marked in bold.
Lotka Volterra SIR
T=1 T=10 T=100 T=1 T=10 T=100
NPE 0.88 0.96 0.99 0.98 0.99 1.00
NPE’ 0.84 0.96 1.00 0.97 0.99 1.00
NPE+ 0.91 0.95 0.99 0.98 0.99 1.00
NPE++ 0.88 0.95 0.99 0.98 1.00 1.00
NLE+ 0.92 0.96 1.00 0.98 0.99 1.00
NLE++ 0.85 0.96 0.99 0.97 0.99 1.00
NSE 0.84 0.95 0.99 0.97 0.99 1.00
FNLE 0.75 0.89 0.98 0.93 0.96 0.96
FNRE 0.83 0.99 0.99 0.94 0.98 1.00
FNSE 0.66 0.90 0.99 0.84 0.89 0.85
NPE 0.85 0.93 0.99 0.96 0.95 1.00
NPE’ 0.83 0.87 0.99 0.98 0.99 1.00
NPE+ 0.87 0.86 0.99 0.96 0.95 1.00
NPE++ 0.87 0.91 0.99 0.96 0.92 1.00
NLE+ 0.87 0.92 0.99 0.96 0.96 1.00
NLE++ 0.84 0.95 0.99 0.96 0.97 1.00
NSE 0.84 0.93 0.99 0.96 0.95 1.00
FNLE 0.60 0.78 0.96 0.93 0.95 0.97
FNRE 0.71 0.98 0.99 0.81 0.89 0.93
FNSE 0.57 0.78 0.94 0.67 0.78 0.74
Table 4: Extended baseline comparison for Lotka Volterra and SIR: Performance per method given as C2ST metric for Lotka Volterra and SIR task. The top half of the table reports results trained on 10k step simulations, and the bottom half on 100k. Best performing methods are marked in bold.
Overall, all “global” methods that aim to reduce the time series into a statically sized summary statistic do behave similarly within our evaluation suite (Table 3, 4). A fundamental challenge with these approaches is sequence length generalization, i.e., finding statistics that generalize to larger lengths than observed in training, which is a challenging task (Ray Chowdhury & Caragea, 2024; Zhang et al., 2022). In contrast, factorized methods do not have this problem. Instead, their main source of error is due to the accumulation of local approximation errors. This can lead to deteriorating performance and is especially visible for NRE within our evaluation. Notably, an increase in C2ST is, to some degree, expected, given that posteriors with more observations likely contract, making slight deviations easier to detect by a classifier. In addition, results are often plotted against the training budget within SBI benchmark (Lueckmann et al., 2017) and not by observation size. We thus visualize our results also in this more similar fashion (see Fig. A6). We performed this on a training budget (equaling the number of transition step evaluations) of $10^{4}$, $10^{5}$, and $10^{6}$ (which corresponds to $10^{3}$, $10^{4}$, and $10^{5}$ “full” simulation for NPE) as in Lueckmann et al. (2021).
Fig 6: Posterior approximation plotted against training budget: Each column corresponds to a task. Each row shows the performance in sliced Wasserstein distance for test observations generated using 1,10, and 100 transitions. In contrast to the main figures, we see the x-axis now specifies the training budget, i.e., the number of transition evaluations that can be used to simulate the training data.
E.2 Calibration analysis Lotka Volterra and SIR
Fig 7: We perform a simulation-based calibration analysis (SBC) (Talts et al., 2018) for the Lotka Volterra and SIR task trained using 100k step simulations using FNSE and NPE. We used 1000 test simulations for evaluation. The shaded Grey area indicates the expected results under the true posterior. Colors indicate the number of transitions of the test simulations.
We performed a simulation-based calibration analysis (Talts et al., 2018) for the Lotka-Volterra and SIR tasks. Consistent with our evaluation using C2ST and sW, we observe that FNSE is better calibrated than NPE across different sequence lengths (Fig. 7ab). This difference is especially pronounced in the SIR task, where NPE completely fails to generalize to sequence lengths larger than those observed during training ($T>10$). In contrast, the effect is less prominent in the Lotka-Volterra task. We hypothesize that, due to the periodic nature of the Lotka-Volterra dynamics, the embedding network more easily identifies representations that generalize well across sequence lengths.
E.3 Composition methods
We compared the results of the different implemented score composition methods. Overall this showed a clear advantage of our implementations of GAUSS/JAC over FNPE (Fig. 10). Even if Langevin corrections are introduced FNPE does tend to perform worse and diverge given a larger number of observations.
@@ E.3 Composition methods
The performance of both the GAUSS and JAC methods is quite similar (Fig. 10, first column), especially in the Gaussian RW task. This is expected as both are equivalent if Gaussian assumptions are not violated. Notably, for samplers with fewer steps on non-Gaussian tasks, JAC tends to perform slightly better than GAUSS on long time series (Fig. 10b, columns two onwards). The imposed regularity conditions on the posterior covariances appear to resolve the numerical challenges associated with this approach, as observed by Linhart et al. (2024).
E.4 Proposal distribution
@@
The choice of proposal distribution can have a significant impact on a finite simulation budget. In our benchmark tasks, we primarily selected simple yet reasonable distributions based on our prior knowledge of the dynamics involved. However, we examine the impact of this choice more closely in the Gaussian RW and Mixture RW tasks (see Fig. 11).
@@
We set the proposal distribution to a Gaussian with a zero mean and variable variance, ranging from very narrow to quite wide. Importantly, at evaluation, we initialize the distribution as a point mass at zero. For a very narrow proposal, we thus anticipate strong performance at $T=1$. However, as $T$ increases, performance may decline because the time series dynamics could extend beyond the high-probability support of the proposal. Conversely, excessively wide proposals may also hinder performance if the values at the observation time fall outside the training set. It is crucial to note that this performance evaluation is relative to our specific evaluation pipeline. In both tasks, evaluations are conducted on random walks that consistently start from zero, thus constraining the range of attainable values given a maximum number of steps.
@@
It is important to emphasize that the proposals we utilized are generally not optimal. Other tasks, such as the Periodic/Linear SDE, Double Well, SIR, and Lotka-Volterra models, exhibit a high correlation among different variables, which the proposal we chose does not capture. See Appendix Sec. 3.4 for some examples. Nevertheless, to ensure a fair comparison with the NPE baseline, we opted for relatively simple proposals.
E.5 Kolmogorov flow
@@
It is important to note that in the NPE approach, we would need to construct an embedding network not only for individual “images” but also for a complete video of arbitrary length, which poses significant challenges.
As an initial distribution, we utilize a filtered velocity field provided by jax_cfd library, with a maximum velocity and peak wave number of three. Our approach closely follows the methodology outlined by Rozet & Louppe (2023). The simulator performs 100 solve steps per transition, with a step size of $10^{-3}$ seconds; each transition thus emulates 10 milliseconds. To introduce stochasticity into the PDE, we add small amounts of Gaussian noise with a standard deviation of $5\cdot 10^{-3}$ after each transition. We used the jax-cfd library (Kochkov et al., 2021) to solve the Navier-Stokes equations on a $64\times 64$ grid.
@@
Notably, this procedure does not violate the requirement that $\tilde{p}({\bm{x}}^{T})$ is independent of ${\bm{\theta}}$, as each transition is performed with a different, independent parameter. This approach is motivated by the fact that, at best, we would like proper samples from the associated prior predictive distribution (Appendix Sec. C).
@@
E.6 Time-Inhomogenous Gaussian Random Work
To evaluate the performance of the time-inhomogeneous extension described in Appendix A.2.1, we consider the Time-Inhomogeneous Gaussian Random Walk (TI Gaussian RW)
with ${\bm{x}}_{t}$ multiplied by the time-inhomogeneous coefficient $\left(\alpha+\frac{1}{t+1}\right)$. We demonstrate the scalability of our methods with dimension $d$ ranging from $1,2$, to $10$, and the number of observation transitions ranging from $1$, $10$, to $100$. As a comparison metric, we use C2ST. The simulation budget is fixed at $10k$ and $100k$, consistent with the problem settings in the main section. Note that the training data include the time variable $t$, and therefore, the training data for estimating the transition distribution $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$ are reduced to $\frac{1}{100}$ compared to the time-homogeneous setting in the main section. The maximum time length $T_{\text{max}}$ in training is set to $100$.
@@ E.6 Time-Inhomogenous Gaussian Random Work
Fig 13: Time inhomogenous Gaussian random walk: We train all methods on 10k (top row) and 100k (bottom row) simulation on the time-inhomogenous Gaussian RW talk. All methods have also been adapted to input the current time (as given by integers ranging from zero to 100).